OpenStudy (anonymous):

Find in the missing step: sec^4 x -2 sec^2 x+1=(sec^2x-1)^2 =????????? =tan^4 x

OpenStudy (anonymous):

(tan^2x)^2

OpenStudy (anonymous):

[ sec^2x-1= tan^2x ]

OpenStudy (anonymous):

how is that?

OpenStudy (anonymous):

is it because tan^2x=sec^2x-1?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

its an identity =)

OpenStudy (anonymous):

okay thanks

OpenStudy (darthsid):

So, the best way to do this is to go back to basics of a right triangle!!  H = hypotenuse, B = base, P = perpendicular \[\sec(x) = \frac{H}{B}\] \[\sec^2(x) = \frac{H^2}{B^2}\] \[Sec^2(x) - 1 = \frac{H^2}{B^2} - 1 = \frac{H^2 - B^2}{B^2}\] Now, according to the Pythagoras theorem, \[H^2 = P^2 + B^2\] Which means \[H^2 - B^2 = P^2\] So, \[Sec^2(x) - 1 = \frac{H^2 - B^2}{B^2} = \frac{P^2}{B^2}\] \[\frac{P}{B} = tan(x)\] Thus, you get your answer :)